40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 8 22:26:03 2016, varun, Update, CDS, DAFI Update: Changes in LSC model lsc.pnglsctodaf.pnglsctodaf1.png
    Reply  Sun Jul 10 20:08:44 2016, varun, Update, CDS, DAFI Update: Changes in LSC and PEM models pemtodaf.pngpemindaf.png
Message ID: 12287     Entry time: Sun Jul 10 20:08:44 2016     In reply to: 12282
Author: varun 
Type: Update 
Category: CDS 
Subject: DAFI Update: Changes in LSC and PEM models 

I have added the PEM signals mentioned in the previous elog as DAF inputs through PCIE IPC, and compiled and restarted the c1pem and c1daf models. 

Attached are the pictures of the simulink diagram of the addition in the PEM and the DAF. 

Since the signals are moving from a 2kHz clock rate machine to a 16kHz clock rate machine, some imaging effects are possible, which I have to look into.

Attachment 1: pemtodaf.png  52 kB  | Hide | Hide all
pemtodaf.png
Attachment 2: pemindaf.png  85 kB  | Hide | Hide all
pemindaf.png
ELOG V3.1.3-